OpenStudy (anonymous):

Differentiate the function. y=4^x + 4/x I guess you start by factoring out the 4 in each term? y= 4(1^x + 1/x) y = 4(1^x +(-1/x^-2)) Hmm... I know that's wrong...

OpenStudy (anonymous):

I got it! d/dx f(x) = d/dx (4^x + 4/x) = d/dx (4^x) + 4 * d/dx (1/x) = 4 + (-d/dx (x) / x^2) + ln 4 * 4^x = ln (4) * 4^x - (4*1/x^2) = ln (4) * 4^x - (4/x^2)